# From paintbrush to pixel: A review of deep neural networks in AI-generated art

### v2

@@ Abstract
MSC:
[…]
journal: Computer Vision and Image Understanding
@@ 1 Introduction
“AI-generated art is a new and emerging field, and it will likely take some time for it to fully mature and for its place in the art world to be fully understood,” replied ChatGPT, the chatbot created by OpenAI, when we asked about the current state of AI-generated art. It continued, stating, “As with any new field, it will take time for it to be fully understood and for its place in the art world to be established.”
@@
Whether they are recognized as art or not, AI-generated images are widespread today. Regardless of discussions about how creative or artistic they are, their existence in our lives cannot be denied any longer. In 2018, the first AI-generated portrait “Edmond de Belamy” (Fig. 1) sold for \$432,500 at Christie’s art auction. It was created using a generative adversarial network (GAN) Goodfellow et al. (2014) as a part of “La Famille de Belamy” series by Obvious Art. The fact that it is signed with the loss function used in GAN makes this case quite amusing. In 2022, Jason M. Allen’s AI-generated artwork, “Théâtre D’opéra Spatial” (Fig. 1), won the art prize in the digital category at the Colorado State Fair’s annual art competition. This piece was created with the AI-based tool Midjourney111www.midjourney.com which can generate images by taking text prompts as input.
@@
Fig. 1: (Left) “Edmond de Belamy” - The first AI-generated portrait sold at Christie’s art auction in 2018. (Right) “Théâtre D’opéra Spatial” - The winner of the digital art category at the Colorado State Fair’s annual art competition in 2022.
@@
The progress in AI-generated art is made possible by deep learning, which is a subfield of machine learning. This subfield includes deep neural networks, which have led to significant breakthroughs in various fields such as computer vision in the last decade. In this paper, we focus on deep neural networks used in image processing and recent developments that have been used to produce AI-generated images. Several studies address AI-generated art from different perspectives. For example, Cetinic et al. (2022) Cetinic and She (2022) discussed the creativity of AI technologies, as well as authorship, copyright and ethical issues. Ragot et al. (2020) conducted an experiment where they asked participants to rate the difference between paintings created by humans and AI in terms of liking, perceived beauty, novelty, and meaning. Other important issues today are credit allocation and responsibility for AI-generated art Epstein et al. (2020). Recently, a review of generative AI models has been released which touches upon various applications, such as texts, images, videos and audios Gozalo-Brizuela and Garrido-Merchan (2023).
@@
In this paper, we focus on the main neural networks which have been used to generate realistic images. We first explain the building blocks of the related neural networks to provide readers with a better understanding of these models. We describe the general working principles of these neural networks, and introduce the recent trends in AI-generated art such as DALL-E 3 Betker et al. (2023). Then, we examine the rapid development of deep neural networks in AI-generated art, and emphasize the milestones in this progress. This review addresses the topic from a technical perspective and provides comparisons of current state-of-the-art models. However, even for a non-technical audience (e.g., from more traditional areas in the fine arts, aesthetics, and cultural studies), this review will provide an overview of the different techniques and tools that are currently available.
@@ 2.1 Convolutional neural networks
Convolutional neural networks (LeCun et al., 1989, 1998), usually referred to as CNNs, or ConvNets, are deep neural networks utilized mainly for image processing. In a deep CNN, increasingly more abstract features are extracted from the image through a series of hidden layers. As such, a CNN follows a similar hierarchy as the human visual cortex, in that earlier areas extract simple visual features and later areas extract combinations of features and high level features. In this manner, the complex mapping of the input pixels is divided into a series of nested mappings, each of which is represented by a different layer Goodfellow et al. (2016).
@@
There are various CNN architectures in the literature (e.g., LeNet LeCun et al. (1989, 1998), which is capable of recognizing handwritten digits). However, more complex tasks like object recognition require deeper CNNs such as AlexNet Krizhevsky et al. (2012), VGG Simonyan and Zisserman (2015), ResNet He et al. (2016), DenseNet Huang et al. (2017), EfficientNet Tan and Le (2019), MobileNet Howard et al. (2017), ResNeXt Xie et al. (2017), Inception and GoogLeNet Szegedy et al. (2015). A typical CNN includes three kinds of layers: convolutional layers, pooling layers, and fully-connected layers. The general structure of a typical CNN for classification is illustrated in Fig. 2. The input image $\mathcal{X}$ is presented at the input layer, which is followed by the first convolutional layer. In a convolutional layer, the weights are kept within kernels. During learning, a mathematical operation called convolution is performed between the input and the kernel. Basically, convolution is performed by multiplying the elements of the input with each element of the kernel and summing the results of this multiplication. This input can be an input image or the output of another, preceding convolutional layer. The units in a convolutional layer are arranged into planes, each of which is referred to as a feature map, defined as 3D tensors. Each unit in a feature map takes input from a small area of the image, and all units detect the same pattern, but at different locations in the input image. All units in a feature map share the same weights, thus a convolution of the image pixel intensities with a kernel is performed Bishop (2006). In general, there are multiple feature maps in a convolutional layer, each with its own weights, to detect multiple features Bishop (2006). Accordingly, a single convolutional layer includes many kernels, each operating at the same time. As a deep neural network, a CNN typically includes many convolutional layers, which implies millions of parameters.
@@
Fig. 2: An example CNN structure with two convolutional, two pooling, and three fully-connected layers for classification.
@@
Each output of the convolution operation is usually run through a nonlinear activation function, such as the rectified linear unit (ReLU) Glorot et al. (2011). Then, a convolutional layer or a stack of convolutional layers is followed by a pooling layer which reduces the size of the feature maps by calculating a summary statistic of the nearby outputs, such as the maximum value or the average Goodfellow et al. (2016). In the end, a series of convolutional and pooling layers is followed by fully-connected (dense) layers. The activation function that the output layer applies depends on the task. In most cases, a sigmoid function is preferred for binary classification, softmax nonlinearity for multi-classification, and linear activation for regression tasks. Accordingly, a CNN minimizes the difference between the desired output values $y$ and the predicted values $\hat{y}$ in the cross-entropy loss functions given in Eq. 1 and Eq. 2 for binary and multi-class classification, respectively. In the case of a regression task, a CNN may minimize mean squared error given in Eq. 3. In the equations below, $W$ represents the weights belonging to the hidden layers, and $V$ represents the output layer weights. Although the aforementioned are the most frequently used activation and loss functions, there are other alternatives available such as Leaky ReLU Maas et al. (2013), SELU Klambauer et al. (2017), Swish Ramachandran et al. (2017), Mish Misra (2020) as activation functions, or mean absolute error as a loss function.
@@
A fully-connected layer and a convolutional layer significantly differ in that fully-connected layers learn global patterns in their input feature space, whereas convolutional layers learn local patterns Chollet (2018). In a CNN, there are fewer weights than there would be if the network were fully-connected because of the local receptive fields Bishop (2006). CNNs are now essential neural networks in deep learning and have yielded major advances for a variety of image processing tasks.
@@ 2.2 Autoencoders
The autoencoder, originally named the autoassociator Cottrell et al. (1987), is a typical example of unsupervised learning. This neural network learns to reconstruct the input data by extracting a (usually lower-dimensional) representation of the data. The autoencoder has been successfully implemented for unsupervised or semi-supervised tasks, or as a preprocessing stage for supervised tasks. The general structure of an autoencoder consists of an encoder and a decoder, as shown in Fig. 3. In the simplest case, both encoder and decoder are composed of fully-connected layer(s). During training, the encoder, in one or more layers, usually transforms the input to a lower-dimensional representation. Then, the decoder that follows, in one or more layers, takes this representation as input and reconstructs the original input back as its output Soydaner (2020). The aim is to obtain a meaningful representation of data, which also makes the autoencoder an example of representation learning.
Fig. 3: The general structure of the (Left) Autoencoder, (Right) Variational autoencoder. $x^{t}$ refers to an input sample, $h^{t}$ to the latent representation and $\hat{x}^{t}$ to the reconstructed input. The parameters of encoder $(\theta_{E})$ and decoder $(\theta_{D})$ are updated during training.
In the general framework, the encoder takes the input $x^{t}$, and produces a compressed or hidden/latent representation of input $h^{t}$. Then, the decoder takes $h^{t}$ as input and reconstructs the original input as the output $\hat{x}^{t}$. When the $h^{t}$ dimension is less than the $x^{t}$ dimension, the autoencoder is called undercomplete. The undercomplete autoencoder can capture the most salient features of the data. On the other hand, when the $h^{t}$ dimension is equal to or greater than the $x^{t}$ dimension, it is called overcomplete. The overcomplete autoencoder just copies the input to the output, as it cannot learn anything meaningful about the data distribution. Regularized autoencoders, such as the sparse autoencoders, alleviate this issue by applying a regularization term in the loss function Goodfellow et al. (2016).
@@
When an autoencoder in which the encoder and decoder are both perceptrons learns a lower-dimensional compressed representation of data, the autoencoder performs similar to principal component analysis (PCA): The encoder weights span the same space spanned by the k highest eigenvectors of the input covariance matrix Bourlard and Kamp (1988). When the encoder and decoder are multilayer perceptrons, then the autoencoder learns to do nonlinear dimensionality reduction in the encoder Soydaner (2020).
Autoencoder architectures extend to specialized forms such as the denoising autoencoder Vincent et al. (2008), which reconstructs clean data from corrupted inputs, and the contractive autoencoder Rifai et al. (2011), which increases robustness to small changes in input data. While discussing the different autoencoder types, we should highlight the variational autoencoder (VAE) Kingma and Welling (2013); Rezende et al. (2014), which turns an autoencoder into a generative model. Similar to the autoencoder architecture mentioned above, a VAE is composed of an encoder and decoder. However, the encoder does not produce a lower-dimensional latent representation. Instead, the encoder produces parameters of a predefined distribution in the latent space for input, i.e., mean and variance. Thus, the input data is encoded as a probability distribution. Then, new samples can be generated by sampling from the latent space that the autoencoder learned during training. There are variants of this neural network architecture, including VQ-VAE Van Den Oord et al. (2017).
@@ 2.3 Generative adversarial networks
The generative adversarial network (GAN) Goodfellow et al. (2014) is a milestone for generative models in deep learning literature. The idea is based on game theory and a minimax game with two players. Instead of human players, in this case, the players are neural networks. One of these neural networks is called a generator (G) while the other is called a discriminator (D). These two networks are trained end-to-end with backpropagation in an adversarial setting, i.e., the generator and discriminator compete with each other. The generator captures the data distribution while the discriminator estimates the probability that its input either comes from the real data or is a fake sample created by the generator (Fig. 4). The competition between these two neural networks in the game makes them improve their results until the fake data generated by G is indistinguishable from the original data for D. As a result, G learns to model the data distribution during training and can generate novel samples after training is completed.
Fig. 4: The general structure of an generative adversarial network (GAN). The generator upscales its input (a noise vector) through a series of layers into an image. The discriminator performs a binary classification task, i.e., deciding whether the input image it receives is real or a generated sample.
In the original GAN framework, both the generator and discriminator are MLPs. The generator takes a random noise vector as input and generates samples. The discriminator takes both the generated sample and a real sample from the data as inputs, and tries to decide which one is real. Then, based on the feedback coming from the discriminator, the generator updates its weights. As shown in the loss function in Eq. 2.3, G tries maximizing the probability of D making a mistake. Ideally, the game ends at a saddle point where D is equal to 1/2 everywhere. In practice, GAN training faces several challenges Khanuja and Agarkar (2023). As a solution, GANs that use different loss functions have been proposed, such as the Wasserstein GAN (WGAN) Arjovsky et al. (2017).
@@
In the literature, there are various GAN architectures Creswell et al. (2018). For example, when the G and D are composed of convolutional layers, this architecture is called a deep convolutional GAN (DCGAN) Radford et al. (2015). StyleGAN Karras et al. (2019) is a convolutional GAN which can vary coarse-to-fine visual features separately. Whereas an ordinary GAN receives a noise vector as input, StyleGAN inputs the noise vector to a mapping network to create an intermediate latent space. The intermediate latents are then fed through the generator through adaptive instance normalization Karras et al. (2019). The mapping network ensures that features are disentangled in latent space, allowing StyleGAN to manipulate specific features in images (e.g., remove someone’s glasses or make someone look older). Pix2pix Isola et al. (2017) is a GAN that takes an image as input (rather than noise) and translates it to a different modality (e.g., BW image to RGB image). Whereas the training of pix2pix requires image pairs (the same image in RGB and BW), CycleGAN Zhu et al. (2017) alleviates this need by ‘cycling’ between two GANs (see Section 3.2). An adversarial autoencoder Makhzani et al. (2015) combines the adversarial setting of GANs with the autoencoder architecture. Lastly, a self-attention adversarial neural network (SAGAN) Zhang et al. (2019) defines a GAN framework with an attention mechanism, which is explained in Section 2.4.

### v3

@@ Abstract
Neural networks, Transformers, Diffusion models, Generative adversarial networks, Text-to-image models, Deep learning, Generative AI, Ethics
@@ 1 Introduction
“AI-generated art is a new and emerging field, and it will likely take some time for it to fully mature and for its place in the art world to be fully understood,” replied ChatGPT [1], the chatbot created by OpenAI, when we asked about the current state of AI-generated art. It continued, stating, “As with any new field, it will take time for it to be fully understood and for its place in the art world to be established.”
@@
Whether they are recognized as art or not, AI-generated images are widespread today. Regardless of discussions about how creative or artistic they are, their existence in our lives cannot be denied any longer. In 2018, the first AI-generated portrait “Edmond de Belamy” (Fig. 1) sold for \$432,500 at Christie’s art auction. It was created using a generative adversarial network (GAN) [2] as a part of “La Famille de Belamy” series by Obvious Art. The fact that it is signed with the loss function used in GAN makes this case quite amusing. In 2022, Jason M. Allen’s AI-generated artwork, “Théâtre D’opéra Spatial” (Fig. 1), won the art prize in the digital category at the Colorado State Fair’s annual art competition. This piece was created with the AI-based tool Midjourney111www.midjourney.com which can generate images by taking text prompts as input.
@@
Figure 1: (Left) “Edmond de Belamy” - The first AI-generated portrait sold at Christie’s art auction in 2018. (Right) “Théâtre D’opéra Spatial” - The winner of the digital art category at the Colorado State Fair’s annual art competition in 2022.
@@
The progress in AI-generated art is made possible by deep learning, which is a subfield of machine learning. This subfield includes deep neural networks, which have led to significant breakthroughs in various fields such as computer vision in the last decade. In this paper, we focus on deep neural networks used in image processing and recent developments that have been used to produce AI-generated images. Several studies address AI-generated art from different perspectives. For example, Cetinic et al. (2022) [3] discussed the creativity of AI technologies, as well as authorship, copyright and ethical issues. [4] conducted an experiment where they asked participants to rate the difference between paintings created by humans and AI in terms of liking, perceived beauty, novelty, and meaning. These are measured through statements such as ‘I like this painting,’ ‘This painting looks beautiful,’ ‘This painting seems novel,’ and ‘I perceive the meaning of the painting.’ They found that identical artworks were evaluated differently depending on whether they were attributed to a human or an AI, and that human-created artworks were generally rated more highly than AI-generated ones. Other important issues today are credit allocation and responsibility for AI-generated art [5]. A review of generative AI models has been released which touches upon various applications, such as texts, images, videos and audios [6]. Recently, Wang et al. (2025) [7] presented a review of diffusion models for visual art creation. In contrast, we provide a broader review of neural networks in image generation for artistic purposes, trace their temporal progress, present accessible technical explanations, and address important issues related to aesthetics and ethics. We include visual comparisons that make the technical discussion more tangible and offer practical guidance for readers selecting a model.
@@
In this paper, we focus on the main neural networks which have been used to generate realistic images. We first explain the building blocks of the related neural networks to provide readers with a better understanding of these models. We describe the general working principles of these neural networks, and introduce the recent trends in AI-generated art such as DALL-E 3 [8]. Then, we examine the rapid development of deep neural networks in AI-generated art, and emphasize the milestones in this progress. This review addresses the topic from a technical perspective and provides comparisons of current state-of-the-art models. However, even for a non-technical audience (e.g., from more traditional areas in the fine arts, aesthetics, and cultural studies), this review will provide an overview of the different techniques and tools that are currently available.
@@ 2.1 Convolutional neural networks
Convolutional neural networks [9, 10], usually referred to as CNNs, or ConvNets, are deep neural networks utilized mainly for image processing. In a deep CNN, increasingly more abstract features are extracted from the image through a series of hidden layers. As such, a CNN follows a similar hierarchy as the human visual cortex, in that earlier areas extract simple visual features and later areas extract combinations of features and high level features. In this manner, the complex mapping of the input pixels is divided into a series of nested mappings, each of which is represented by a different layer [11].
@@
There are various CNN architectures in the literature (e.g., LeNet [9, 10], which is capable of recognizing handwritten digits). However, more complex tasks like object recognition require deeper CNNs such as AlexNet [12], VGG [13], ResNet [14], DenseNet [15], EfficientNet [16], MobileNet [17], ResNeXt [18], Inception and GoogLeNet [19]. A typical CNN includes three kinds of layers: convolutional layers, pooling layers, and fully-connected layers. The general structure of a typical CNN for classification is illustrated in Fig. 2. The input image $\mathcal{X}$ is presented at the input layer, which is followed by the first convolutional layer. In a convolutional layer, the weights are kept within kernels. During learning, a mathematical operation called convolution is performed between the input and the kernel. Basically, convolution is performed by multiplying the elements of the input with each element of the kernel and summing the results of this multiplication. This input can be an input image or the output of another, preceding convolutional layer. The units in a convolutional layer are arranged into planes, each of which is referred to as a feature map, defined as 3D tensors. Each unit in a feature map takes input from a small area of the image, and all units detect the same pattern, but at different locations in the input image. All units in a feature map share the same weights, thus a convolution of the image pixel intensities with a kernel is performed [20]. In general, there are multiple feature maps in a convolutional layer, each with its own weights, to detect multiple features [20]. Accordingly, a single convolutional layer includes many kernels, each operating at the same time. As a deep neural network, a CNN typically includes many convolutional layers, which implies millions of parameters.
@@
Figure 2: An example CNN structure with two convolutional, two pooling, and three fully-connected layers for classification.
@@
Each output of the convolution operation is usually run through a nonlinear activation function, such as the rectified linear unit (ReLU) [21]. Then, a convolutional layer or a stack of convolutional layers is followed by a pooling layer which reduces the size of the feature maps by calculating a summary statistic of the nearby outputs, such as the maximum value or the average [11]. In the end, a series of convolutional and pooling layers is followed by fully-connected (dense) layers. The activation function that the output layer applies depends on the task. In most cases, a sigmoid function is preferred for binary classification, softmax nonlinearity for multi-classification, and linear activation for regression tasks. Accordingly, a CNN minimizes the difference between the desired output values $y$ and the predicted values $\hat{y}$ in the cross-entropy loss functions given in Eq. 1 and Eq. 2 for binary and multi-class classification, respectively. In the case of a regression task, a CNN may minimize mean squared error given in Eq. 3. In the equations below, $W$ represents the weights belonging to the hidden layers, and $V$ represents the output layer weights. Although the aforementioned are the most frequently used activation and loss functions, there are other alternatives available such as Leaky ReLU [22], SELU [23], Swish [24], Mish [25] as activation functions, or mean absolute error as a loss function.
@@
A fully-connected layer and a convolutional layer significantly differ in that fully-connected layers learn global patterns in their input feature space, whereas convolutional layers learn local patterns [26]. In a CNN, there are fewer weights than there would be if the network were fully-connected because of the local receptive fields [20]. CNNs are now essential neural networks in deep learning and have yielded major advances for a variety of image processing tasks.
@@ 2.2 Autoencoders
The autoencoder, originally named the autoassociator [27], is a typical example of unsupervised learning. This neural network learns to reconstruct the input data by extracting a (usually lower-dimensional) representation of the data. The autoencoder has been successfully implemented for unsupervised or semi-supervised tasks, or as a preprocessing stage for supervised tasks. The general structure of an autoencoder consists of an encoder and a decoder, as shown in Fig. 3. In the simplest case, both encoder and decoder are composed of fully-connected layer(s). During training, the encoder, in one or more layers, usually transforms the input to a lower-dimensional representation. Then, the decoder that follows, in one or more layers, takes this representation as input and reconstructs the original input back as its output [28]. The aim is to obtain a meaningful representation of data, which also makes the autoencoder an example of representation learning.
Figure 3: The general structure of the (Left) Autoencoder, (Right) Variational autoencoder. $x^{t}$ refers to an input sample, $h^{t}$ to the latent representation and $\hat{x}^{t}$ to the reconstructed input. The parameters of encoder $(\theta_{E})$ and decoder $(\theta_{D})$ are updated during training.
In the general framework, the encoder takes the input $x^{t}$, and produces a compressed or hidden/latent representation of input $h^{t}$. Then, the decoder takes $h^{t}$ as input and reconstructs the original input as the output $\hat{x}^{t}$. When the $h^{t}$ dimension is less than the $x^{t}$ dimension, the autoencoder is called undercomplete. The undercomplete autoencoder can capture the most salient features of the data. On the other hand, when the $h^{t}$ dimension is equal to or greater than the $x^{t}$ dimension, it is called overcomplete. The overcomplete autoencoder just copies the input to the output, as it cannot learn anything meaningful about the data distribution. Regularized autoencoders, such as the sparse autoencoders, alleviate this issue by applying a regularization term in the loss function [11].
@@
When an autoencoder in which the encoder and decoder are both perceptrons learns a lower-dimensional compressed representation of data, the autoencoder performs similar to principal component analysis (PCA): The encoder weights span the same space spanned by the k highest eigenvectors of the input covariance matrix [29]. When the encoder and decoder are multilayer perceptrons, then the autoencoder learns to do nonlinear dimensionality reduction in the encoder [28].
Autoencoder architectures extend to specialized forms such as the denoising autoencoder [30], which reconstructs clean data from corrupted inputs, and the contractive autoencoder [31], which increases robustness to small changes in input data. While discussing the different autoencoder types, we should highlight the variational autoencoder (VAE) [32, 33], which turns an autoencoder into a generative model. Similar to the autoencoder architecture mentioned above, a VAE is composed of an encoder and decoder. However, the encoder does not produce a lower-dimensional latent representation. Instead, the encoder produces parameters of a predefined distribution in the latent space for input, i.e., mean and variance. Thus, the input data is encoded as a probability distribution. Then, new samples can be generated by sampling from the latent space that the autoencoder learned during training. There are variants of this neural network architecture, including VQ-VAE [34].
@@ 2.3 Generative adversarial networks
The generative adversarial network (GAN) [2] is a milestone for generative models in deep learning literature. The idea is based on game theory and a minimax game with two players. Instead of human players, in this case, the players are neural networks. One of these neural networks is called a generator (G) while the other is called a discriminator (D). These two networks are trained end-to-end with backpropagation in an adversarial setting, i.e., the generator and discriminator compete with each other. The generator captures the data distribution while the discriminator estimates the probability that its input either comes from the real data or is a fake sample created by the generator (Fig. 4). The competition between these two neural networks in the game makes them improve their results until the fake data generated by G is indistinguishable from the original data for D. As a result, G learns to model the data distribution during training and can generate novel samples after training is completed.
Figure 4: The general structure of a generative adversarial network (GAN). The generator upscales its input (a noise vector) through a series of layers into an image. The discriminator performs a binary classification task, i.e., deciding whether the input image it receives is real or a generated sample.
In the original GAN framework, both the generator and discriminator are MLPs. The generator takes a random noise vector as input and generates samples. The discriminator takes both the generated sample and a real sample from the data as inputs, and tries to decide which one is real. Then, based on the feedback coming from the discriminator, the generator updates its weights. As shown in the loss function in Eq. 2.3, G tries maximizing the probability of D making a mistake. Ideally, the game ends at a saddle point where D is equal to 1/2 everywhere. In practice, GAN training faces several challenges [35]. As a solution, GANs that use different loss functions have been proposed, such as the Wasserstein GAN (WGAN) [36].
@@
In the literature, there are various GAN architectures [37]. For example, when the G and D are composed of convolutional layers, this architecture is called a deep convolutional GAN (DCGAN) [38]. StyleGAN [39] is a convolutional GAN which can vary coarse-to-fine visual features separately. Whereas an ordinary GAN receives a noise vector as input, StyleGAN inputs the noise vector to a mapping network to create an intermediate latent space. The intermediate latents are then fed through the generator through adaptive instance normalization [39]. The mapping network ensures that features are disentangled in latent space, allowing StyleGAN to manipulate specific features in images (e.g., remove someone’s glasses or make someone look older). Pix2pix [40] is a GAN that takes an image as input (rather than noise) and translates it to a different modality (e.g., BW image to RGB image). Whereas the training of pix2pix requires image pairs (the same image in RGB and BW), CycleGAN [41] alleviates this need by ‘cycling’ between two GANs (see Section 3.2). An adversarial autoencoder [42] combines the adversarial setting of GANs with the autoencoder architecture. Lastly, a self-attention adversarial neural network (SAGAN) [43] defines a GAN framework with an attention mechanism, which is explained in Section 2.4.